OpenStudy (anonymous):

A diner has collected data about customer coffee-drinking habits. They have calculated that P(cream) = 0.5, P(sugar) = 0.6, and P(cream or sugar) = 0.7. Determine the P(cream and sugar).

OpenStudy (amistre64):

P(a or b) = P(a) + P(b) - P(a and b)

OpenStudy (amistre64):



OpenStudy (amistre64):

P(a) = x+y P(b) = y+z P(a and b) = y P(a or b) = x + y + z a+b = x+y+y+z = x+2y+z, so we have to subtract a 'y' to adjust